# TrendFact: A Benchmark for Explainable Hotspot Perception in Fact-Checking with Natural Language Explanation

### v3

@@ Abstract
Although fact verification remains fundamental, explanation generation serves as a critical enabler for trustworthy fact-checking systems by producing interpretable rationales and facilitating comprehensive verification processes. However, current benchmarks have limitations that include the lack of impact assessment, insufficient high-quality explanatory annotations, and an English-centric bias. To address these, we introduce TrendFact, the first hotspot perception fact-checking benchmark that comprehensively evaluates fact verification, evidence retrieval, and explanation generation tasks. TrendFact consists of 7,643 carefully curated samples sourced from trending platforms and professional fact-checking datasets, as well as an evidence library of 66,217 entries with publication dates. We further propose two metrics, ECS and HCPI, to complement existing benchmarks by evaluating the system’s explanation consistency and hotspot perception capability, respectively. Experimental results show that current fact-checking systems, including advanced RLMs such as DeepSeek-R1, face significant limitations when evaluated on TrendFact, highlighting the real-world challenges posed by it. To enhance the fact-checking capabilities of reasoning large language models (RLMs), we propose FactISR, which integrates dynamic evidence augmentation, evidence triangulation, and an iterative self-reflection mechanism. Accordingly, FactISR effectively improves RLM performance, offering new insights for explainable and complex fact-checking.
@@ 1 Introduction
The proliferation of counterfeit claims poses significant societal risks, including mass panic, social destabilization, and even armed conflicts, as exemplified by the COVID-19 infodemic van Der Linden et al. (2020); Aondover et al. (2024). This critical challenge has driven substantial research efforts in automated fact-checking systems, particularly in developing comprehensive benchmark datasets. The rapid expansion of open datasets has accelerated advancements in AI-powered verification technologies, especially through large language models (LLMs) Atanasova (2024); Rani et al. (2023); Wang and Shu (2023); Bilal et al. (2024); Kao and Yen (2024). Despite these developments, current fact-checking benchmarks exhibit several critical limitations:
@@
First, as an emerging subtask in the field of fact-checking, explanation generation plays a pivotal role in enhancing the interpretability of verification results and refining the fact-checking process. However, existing benchmarks mostly lack textual explanations, making them inadequate to effectively evaluate the explanation generation performance. This limitation not only hinders the comprehensive evaluation of fact-checking but also restricts the improvement of the interpretability.
Second, with rapidly emerging information, the ability of fact-checking systems to perceive and check hotspot events has become increasingly important. However, existing benchmarks mostly focus on creating diverse and challenging samples while neglecting the hotspot characteristics of events. This renders them incapable of effectively measuring fact-checking systems’ ability to respond to hotspot events. Moreover, the absence of Chinese benchmarks also hampers the comprehensive development of AI in the field of fact-checking.
To that end, we introduce TrendFact, the first hotspot perception fact-checking benchmark in Chinese scenario that incorporates a comprehensive evaluation of fact verification, evidence retrieval, and explanation generation tasks. It contains 7,643 samples collected from several trending platforms and existing datasets, as well as an evidence library of over 66,217 entries. Each sample in TrendFact consists of attributes including: claim, label, gold evidence, textual explanation, influence score, and four hotspot indicators. TrendFact dataset covers five domains such as healthcare and economy, as well as various reasoning types like numerical calculation and logical inference. Additionally, we propose two new metrics, ECS (Explanation Consistency Score) and HCPI (Hotspot Claim Perception Index), to support the evaluation of fact-checking systems on explanation generation and hotspot perception capability on TrendFact.
@@
The construction of TrendFact consists of three components: (1) Data Collection; (2) Data Augmentation; (3) Evidence Library Construction. For the data collection, we collect a large-scale real-world fact entries from Weibo, Baidu and DouYin and existing dataset CHEF as initial data. For the data augmentation, we conduct two separate augmentation process for the initial data, including rigorous filtering process, claim rewrite, gold evidence retrieval, and textual explanation annotate. For the evidence library construction, we combine the gold evidence from dataset and the evidence retrieved from specifically defined process.
@@
Additionally, We propose FactISR (Augmenting Fact-Checking via Iterative Self-Reflection), a methodology that enhances RLMs on fact-checking tasks by combining dynamic evidence augmentation, reasoning adaptation through evidence triangulation, and iterative self-reflection via reward decoding. We evaluate TrendFact on advanced LLMs, RLMs and fact-checking methods. The experiments results show that most test methods exhibit limited performance and our method FactISR introduce a improvement when applied to RLMs.
@@
  • We introduce TrendFact, a hotspot perception fact-checking benchmark in Chinese scenario that incorporates comprehensive evaluation on evidence retrieval, fact verification, and explanation generation tasks. To the best of our knowledge, it is the first benchmark that address the challenge of hostspot perception and explanation generation capability evaluation of fact-checking systems.
  • We propose FactISR, which integrates dynamic evidence addition, evidence triangulation, and an iterative self-reflection mechanism to enhance the reasoning ability of RLMs on fact-checking task.
  • We propose two new metrics ECS and HCPI to evaluate fact-checking systems on explanation generation and hotspot perception.
  • We conduct extensive experiments that reveal limitations in the capabilities of LLMs, RLMs and fact-checking methods on TrendFact. Additionally, the results show that FactISR can enhance the performance of RLMs [^1] Code and data are available at https://github.com/
@@
Figure 1: Overview of TrendFact. The left side illustrates the diverse data sources of TrendFact, the detailed distribution of the data, and the process of the benchmark construction. The right side displays a fact-checking example from TrendFact that involves numerical reasoning.
2 Related Work
Fact-checking Benchmarks
Existing fact-checking benchmarks can be divided into two primary categories based on their data sources. The first category includes benchmarks derived from Wikipedia data, such as STATPROPS Thorne and Vlachos (2017), FEVEROUS Aly et al. (2021), and Hover Jiang et al. (2020). The second category focuses on datasets developed by refining knowledge bases from fact-checking websites and existing fact-checkers, such as CLAIMDECOMPChen et al. (2022), DECLAREPopat et al. (2018), and QUANTEMPVenktesh et al. (2024). These benchmarks primarily focus on crafting datasets of diverse data types and challenging reasoning tasks and mainly target fact verification and evidence retrieval task, while overlooking the assessment of explanation generation. As LLMs play an increasing role in fact-checking, particularly by generating explanations to aid in verification, it is crucial to assess the reliability of this process. Additionally, evaluating the ability of fact-checking systems to perceive hotspot events is also a pressing issue in the field. However, most existing benchmarks neglect the two key needs above. Moreover, only a few benchmarksGupta and Srikumar (2021); Lin et al. (2024); Hu et al. (2022) have paid attention to datasets in non-English contexts. Therefore, to comprehensively evaluate fact-checking systems, it is essential to develop a non-English benchmark that includes explanations and can assess the system’s ability to perceive hotspots.
Automatic Fact-checking
Research on automated fact-checking primarily falls into two categories: fact verification and explanation generation. Fact verification focuses on timely claim evaluation and has been widely explored in contexts such as Wikipedia articles, table-based data, and QA dialogues. With the rise of LLMs, methods like PROGRAMFC Pan et al. (2023) generate executable programs to support step-by-step verification. Explanation generation aims to produce interpretable outputs, but most work treats it as an intermediate step rather than a core objective. Few studies explore using natural language to convey both claim veracity and reasoning, which is critical for model interpretability and user understanding. For instance, He et al. (2023) generate counter-misinformation responses to correct false claims. Overall, current approaches often neglect the interplay between verification and explanation, limiting both transparency and user trust.
Dataset #Claims Source Language Explanation Focus Hotspot Perception Task
@@ 2 Related Work
Table 1: Comparison of TrendFact with other fact-checking datasets. In the table, WP refers to Wikipedia, FCS refers to fact-checking websites, TP refers to Trending Platform, and FB refers to FreeBase. By “Explanation Focus”, here we refer to dataset contains explanations.
@@ 3 TrendFact Construction
In this section, we describe the fact-checking benchmark TrendFact which comprises 7,643 fact-checking samples along with an evidence library of 66,217 evidences. Figure 1 provides an overview of the dataset attributes, the benchmark construction process, and an sample from TrendFact. Next, we will introduce the data attributes and the construction process of TrendFact in detail.
@@ 3.1 Data Attributes
The TrendFact includes data in five domains, such as public health, science, society, politics and culture. These data are sourced from Trending Platforms (Weibo, Baidu, DouYin) and existing fact-checking dataset CHEF. Each sample in TrendFact consists of attributes including: claim, label, gold evidence, textual explanation, influence score, and four hotspot indicators. Depending on the number of gold evidence items, the dataset is divided into two categories: single-evidence and multi-evidence. Among these, samples with a single gold evidence account for 85%. For more details on hotspot indicators, importance scores, and other attributes, see Appendix A.
3.2 Benchmark Construction
We divide the construction process of TrendFact into three modules: data collection, data augmentation, and evidence library construction. Each module will be described in detail below.
3.2.1 Data Collection
In contrast to previous datasets, which primarily relied on data sources with limited timeliness such as Wikipedia and fact-checking websites, we select more practical and up-to-date sources that can capture emerging trends in a timely manner, including:
  • Trending Platforms. Trending platforms such as Weibo, DouYin, and Baidu contain a wealth of dynamic and up-to-date factual content that can be leveraged for fact-checking. For example, their daily hotspot ranking pages often include concise factual statements accompanied by hotspot indicators such as view counts and discussion volumes. These indicators carry real-time dynamic features that are crucial for evaluating whether a fact-checking system possesses hotspot perception capabilities. Moreover, the content filtering mechanisms of trending platforms provide a certain level of quality assurance for the factual information they present. Therefore, we collect approximately 500,000 hotspot event entries from these platforms between 2020 and 2024 as part of the initial dataset.
  • Existing Datasets. Although the data sources of existing fact-checking datasets have dynamic limitations, they still offer valuable information for broadening verification scenarios beyond trending platforms. However, only a few datasets specifically focus on the Chinese scenario. For example, the CHEF dataset, which collects data from multiple fact-checking websites can serve as a representative resource. Therefore, we include CHEF as part of the data sources in our initial dataset.
Figure 2: Overview of FactISR. The bottom section depicts the token generation process, where RLMs’ thought is enclosed by <think> and </think>. When the "Symbol" token is hit, it provides a reward for the next token being a "Reflection" token. The topmost section shows DEA’s dynamic evidence addition, while the middle-left details ET’s reasoning steps. The middle-right illustrates the decreasing reward for "Reflection" tokens with increased reflections.
3.2.2 Data Augmentation
The initial data collected from trending platforms and existing fact-checking datasets exhibit distinct limitations. First, the trending data lacks the key attributes required for serve as effective fact-checking samples and mostly consists of meaningless noise. Second, the evidence in the CHEF dataset often includes judgmental labels (e.g., “… is a rumor”), which reduces the difficulty of the fact-checking task. In addition, CHEF also contains noisy samples with factual inaccuracies. More information are detailed in Appendix B. Lastly, both data sources lack textual explanations, which are essential for TrendFact. To address these issues, we recruit an annotation team consisting of fourteen graduate students and one doctoral student to carry out a series of rigorous data augmentation processes for each of the two intial dataset.
Data from Trending Platforms
We first employ a voting mechanism with multiple LLMs to filter out 90% of the meaningless samples and those lacking sufficient challenge potential. Subsequently, our annotation team further screens the remaining samples to remove sensitive or overlapping data, resulting in a final set of 6,512 samples. Since the initial trending data lacked the essential factual elements required for fact-checking claims, these samples could not be directly used as claims. To address this issue, we develop a rewriting guideline, and annotators manually rewrite the 6,512 samples accordingly. Consequently, for each of rewritten sample, we then manually retrieve one piece of gold evidence and provide a detailed explanatory annotation (more details in Appendix C).
Data from Existing Datasets
For the CHEF dataset, we first manually filter out noisy data containing sensitive content and correct samples with factual errors. Then, similar to the trending data filtering, we apply an LLM voting mechanism to select samples that present significant verification challenges. Each of the remaining samples is then annotated with a detailed explanation by our annotation team. Consequently, the enhanced CHEF dataset comprises 1,131 high-quality samples.
After augmentation, we merge the two augmented datasets to construct the TrendFact dataset, which contains a total of 7,643 samples. Notably, we provide rigorous training to our annotation team and conduct standardized human evaluations on both the rewritten and annotated results to ensure data quality. The evaluation criteria are detailed in Appendix D. While TrendFact primarily focuses evaluating the hotspot perception capabilities of fact-checking systems, we also aim to extend its utility to assess fundamental verification abilities. To this end, we set the hotspot indicator attribute to null for approximately 15% of the samples.
3.2.3 Evidence Library Construction
@@
The Evidence Library is a crucial component of the fact-checking benchmark, providing essential support for accurate verification. Specifically, we first employ an LLM to generate a set of candidate claims that are semantically similar to the original claim but differ in content and structure. We then retrieve supporting evidence for these candidates from websites and combine it with the gold evidence from the TrendFact dataset to construct a comprehensive evidence library comprising 66,217 entries. The evidence library construction process is simply illustrated in Figure 1 and more detailed descriptions of the evidence library construction process is provided in Appendix E.
3.3 Comparisons with Existing Benchmarks
We conduct a comprehensive comparative analysis between TrendFact and existing fact-checking benchmarks, with the results summarized in Table 1. Our analysis reveals that most existing benchmarks fall short in evaluating explanation quality and hotspot perception capabilities, the two critical aspects for advancing the fact-checking field. These limitations significantly hinder the development of more robust and responsive verification systems. In contrast, TrendFact offers a comprehensive evaluation across three key fact-checking tasks and effectively assesses a system’s ability to respond to emerging hotspot events.

### v4

@@ Abstract
Fact-checking benchmarks provide standardized testing criteria for automated fact-checking systems, driving technological advancement. With the surge of misinformation on social media and the emergence of various fact-checking methods, public concern about the transparency of automated systems and the accuracy of fact-checking for high infulence events has grown. However, existing benchmarks fail to meet these urgent needs and are predominantly English-centric, hindering the progress of comprehensive fact-checking. To address these issues, we introduce TrendFact, the first benchmark capable of evaluating hotspot perception ability (HPA) and all fact-checking tasks. TrendFact consists of 7,643 curated samples sourced from trending platforms and professional fact-checking datasets, as well as an evidence library containing 366,634 entries with publication dates. Additionally, to complement existing benchmarks in evaluating system explanation consistency and HPA, we propose two new metrics: ECS and HCPI. Experimental results show that current fact-checking systems face significant limitations when evaluated on TrendFact, which facilitates the development of more robust fact-checking methods. Furthermore, to enhance the capabilities of existing advanced fact-checking systems, the reasoning large language models (RLMs), we propose FactISR, a reasoning framework that integrates dynamic evidence augmentation with influence score-based iterative self-reflection. FactISR effectively improves RLM’s performance, offering new insights into explainable and complex fact-checking.
@@ 1 Introduction
Figure 1: A fact-checking example from TrendFact that involves numerical reasoning.
@@
The proliferation of counterfeit claims poses significant social risks, including mass panic, social destabilization, and even armed conflicts, as exemplified by the COVID-19 infodemic van Der Linden et al. (2020); Aondover et al. (2024). This critical challenge drives substantial research efforts in automated fact-checking systems, particularly in developing comprehensive fact-checking benchmarks. The rapid expansion of open datasets has accelerated advancements in AI-powered verification technologies, especially through large language models (LLMs) and reasoning large language models (RLMs) Atanasova (2024); Rani et al. (2023); Wang and Shu (2023); Bilal et al. (2024); Kao and Yen (2024). Despite these developments, current benchmarks exhibit two key limitations.
@@
First, as an emerging subtask in the field of fact-checking, explanation generation plays a pivotal role in enhancing the interpretability of verification results and refining the fact-checking process. However, existing benchmarks mostly lack textual explanations, making them inadequate to effectively evaluate the explanation generation performance. As our overview in Table 1 illustrates, this limitation not only hinders the comprehensive evaluation of fact-checking but also restricts the improvement of interpretability.
Second, different claims have varying degrees of social influence and should receive different levels of attention from fact-checking systems. And the influence typically depends on two key factors: popularity and type. However, due to the lack of social hotspot indicators and category classification, existing benchmarks evaluate fact-checking systems equally across all claims, failing to assess the system’s performance in handling high-influence events. Moreover, most existing datasets focus on English-language contexts, and the lack of Chinese benchmarks hinders the comprehensive development of AI in fact-checking.
We define the fact-checking system’s Hotspot Perception Ability (HPA) as its effectiveness in handling high-influence events and propose TrendFact, the first benchmark capable of evaluating HPA and covering all fact-checking tasks. TrendFact focuses on the Chinese language, contains 7,643 samples with an evidence library of 366,634 pieces, and supports all fact-checking tasks, including fact verification, evidence retrieval, and explanation generation. Each sample includes 10 attributes, as shown in Figure 1. Among them, the influence score is derived from the risk score and hotspot indicators. Furthermore, TrendFact evaluates HPA through our proposed metric, HCPI (Hotspot Claim Perception Index), and introduces the ECS (Explanation Consistency Score) metric to fill the gap in evaluating explanation generation consistency.
@@
The construction of the TrendFact benchmark consists of four components: (1) Data Collection, (2) Data Augmentation, (3) Evidence Library Construction, and (4) ECS and HCPI Metrics Definition. First, we gather large-scale real-world fact entries from sources like Weibo, Baidu, etc., along with the existing CHEF dataset as the initial data. Then, we apply two separate augmentation processes to the two data sources. For the construction of evidence library, we combine gold evidence from the dataset with evidence retrieved through our defined process. Finally, we provide formal definitions for the ECS and HCPI metrics.
@@
Furthermore, to address the performance degradation of RLMs caused by the predefined fixed retrieval in traditional retrieve augmentation generation (RAG), as discussed in previous work Li et al. (2025), we propose FactISR, a reasoning framework that enhances RLMs’ performance and HPA on fact-checking tasks by combining dynamic evidence augmentation with influence score-based iterative self-reflection. We evaluate TrendFact on advanced LLMs, RLMs and fact-checking methods. The experiment results show that most tested methods exhibit limited performance, and our FactISR achieves performance improvements with improved computational efficiency.
@@
  • We first define the fact-checking system’s Hotspot Perception Ability (HPA) as its effectiveness in handling high-influence events, and introduce the first Chinese-language benchmark that provides comprehensive evaluations on evidence retrieval, fact verification, explanation generation and HPA, along with two new metrics: ECS for evaluating explanation generation consistency and HCPI for assessing HPA.
  • We propose FactISR, a reasoning framework that enhances RLMs on fact-checking tasks by combining dynamic evidence augmentation with influence score-based iterative self-reflection.
  • We conduct extensive experiments that reveal limitations in the capabilities of LLMs, RLMs, and fact-checking methods on TrendFact. Furthermore, the results show that FactISR can enhance the performance of RLMs [^1] Code and data are available at https://github.com/
@@
Dataset #Claims Source Language Explanation Contains HPA Task
@@
Table 1: Comparison of TrendFact with other fact-checking datasets. WP refers to Wikipedia, FCS refers to fact-checking websites, FB refers to FreeBase, and TP refers to Trending Platforms. By ’HPA’, we refer to whether the dataset supports the evaluation of a fact-checking system’s hotspot perception ability.
2 Related Work
Fact-checking Benchmarks
Existing fact-checking benchmarks are generally divided into two categories: one based on Wikipedia data, such as Hover Jiang et al. (2020) and FEVER Thorne et al. (2018), and another using knowledge bases from fact-checking websites, such as CLAIMDECOMP Chen et al. (2022) and QUANTEMP Venktesh et al. (2024) (Table 1). These benchmarks primarily focus on fact verification and evidence retrieval tasks, often neglecting explanation generation. With the growing role of LLMs and RLMs in generating explanations, evaluating their reliability is crucial. Moreover, it is also a pressing issue in the field to evaluate the HPA of fact-checking systems Solovev and Pröllochs (2022); Sehat et al. (2024), which refers to the system’s greater effort in addressing high influence events. Few benchmarks Gupta and Srikumar (2021); Lin et al. (2024); Hu et al. (2022) have addressed datasets in non-English languages. Therefore, developing a Chinese benchmark like TrendFact, which evaluates both explanation generation and hotspot perception, is essential for comprehensive fact-checking evaluation.
Automatic Fact-checking
Research on automated fact-checking primarily falls into two categories: fact verification and explanation generation. Fact verification focuses on timely claim evaluation and has been widely explored in contexts such as Wikipedia articles, table-based data, and QA dialogues. With the rise of LLMs and RLMs, methods like PROGRAMFC Pan et al. (2023) generate executable programs to support step-by-step verification. Explanation generation aims to produce interpretable outputs, but most work treats it as an intermediate step rather than a core objective. Few studies explore using natural language to convey both claim veracity and reasoning, which is critical for model interpretability and user understanding. For instance, He et al. (2023) generates counter-misinformation responses to correct false claims. Additionally, many methods overlook the HPA, which is essential for improving fact-checking systems’ ability to solve highly influential claims. Incorporating HPA, along with the interplay between verification and explanation, can enhance transparency, trust, and the solving capability of high influence claims in the real world.
Figure 2: The overall construction process of TrendFact includes claim collection, filtering, augmentation, evidence library construction, and a multi-stage sample review process.
3 HotSpot Perception Ability (HPA)
With the explosive growth of information and the accelerated spread on social media, fact-checking systems face significant challenges and need to account for the variations in social influence across events Solovev and Pröllochs (2022); Sehat et al. (2024); Damasceno et al. (2024). For high-influence events, such as misinformation in politics and public health, the system should prioritize these events and allocate more resources. To this end, we introduce the Hotspot Perception Ability (HPA) to measure a fact-checking system’s capacity under such conditions, which we formally define as:
$\text{HPA}=f(\text{C},\text{HI},\text{R})$ (1)
Where C means event category, HI refers to the Hotspot Indicators of an event (e.g., popularity, discussion volume), and R represents the computational resources. By enhancing HPA, the system can dynamically adjust its resource allocation, improving the effectiveness and efficiency of handling events with varying influence.
@@ 4 TrendFact Construction
In this section, we detail the overall construction process of TrendFact, as illustrated in Figure 2. We begin by collecting claims from two different sources (Section 4.2) and applying two separate augmentation processes to refine the collected claims (Section 4.3). Then, we construct an evidence library (Section 4.4). Finally, to enable TrendFact to evaluate both the HPA and the explanation generation consistency of fact-checking systems, we introduce two new metrics: ECS and HCPI (Section 4.5).
@@ 4.1 Data Attributes
TrendFact covers five domains: public health, science, society, politics, and culture. As shown in Figure 1, each sample is annotated with attributes including claim, label, gold evidence, textual explanation, risk score, influence score, and four hotspot indicators. Based on the number of gold evidence items, samples are categorized into single-evidence and multi-evidence, with single-evidence samples accounting for 85%. For more detailed definitions of hotspot indicators, influence scores, and other attributes, see Appendix A.
4.2 Claim Collection
In contrast to previous datasets that primarily relied on sources with limited timeliness and popularity, such as Wikipedia and fact-checking websites, TrendFact collects claims from two complementary sources. Source One: Trending Platforms. We collect claims from Weibo, DouYin, and Baidu, which provide large-scale, dynamic factual statements accompanied by real-time hotspot indicators (e.g., view counts and discussion volumes). These features are crucial for evaluating the HPA of fact-checking systems. In total, we obtain approximately 500,000 claims from these platforms between 2020 and 2024. Source Two: Existing Datasets. We further collect claims from existing fact-checking datasets, which broaden verification scenarios beyond trending platforms. Although such datasets lack hotspot indicators, they still offer valuable resources. In particular, we incorporate claims from the CHEF dataset, which aggregates data from multiple fact-checking websites, as a representative supplement.
4.3 Data Augmentation
The initial claims collected from trending platforms and existing datasets exhibit distinct limitations. Claims from trending platforms often lack the key attributes required for fact-checking (e.g., labels, evidence, and explanations) and contain a large proportion of meaningless noise. In contrast, claims from the existing dataset CHEF, while annotated with labels and evidence, frequently include judgmental words (e.g., “… is a rumor”), noisy samples, and lack textual explanations. To address these issues, we design tailored augmentation processes for the two claim sources.
Trending Claims.
We first apply an LLM voting mechanism to filter out approximately 90% of low-quality or trivial claims (the filtering principle is illustrated in Appendix Table 8). Then, our annotation team removes sensitive or overlapping content, resulting in a set of 6,512 claims. Since the remaining claims often lack the essential factual elements and cannot be used directly as fact-checking samples, annotators follow a rewriting guideline to produce qualified claims. This rewriting process adheres to five fact-checking claim rewriting factors, including: Temporal Anchoring: Adding or specifying temporal references. Data Granularity: Breaking down aggregated data into verifiable units. Ambiguity Resolution: Replacing fuzzy qualifiers with deterministic statements. Comparative Standard: Introducing quantifiable comparison objects and proportions. Domain Knowledge: Supplementing with industry-specific parameters or mechanisms. More details are in Appendix C.
For each rewritten claim, annotators then retrieve one piece of gold evidence and provide a detailed manual explanation. More examples of filtering and rewriting are shown in Appendix Figure 7 and Figure 8, respectively.
CHEF Claims.
We first remove sensitive or factually inaccurate samples and then use the same LLM voting mechanism to select claims that present significant verification challenges. Each selected claim is further annotated with a detailed explanation by our annotation team that shares the same standard applied to trending claims. This process results in 1,131 enhanced claims.
After augmentation, we merge the two sets of claims to construct TrendFact, which consists of 7,643 samples. To ensure the quality of the augmented data, annotators underwent rigorous training, including understanding and applying the filtering mechanisms, rewriting guidelines, and reviewing the claim annotations. The training process also covered examples of proper claim rewriting, emphasizing the importance of maintaining semantic integrity while transforming the claim into a verifiable, fact-checkable statement. As shown in Figure 2, all rewritten and annotated outputs are validated through a three-level standardized human evaluation criteria, ensuring consistency and reliability across the dataset (see Appendix D for details).
4.4 Evidence Library Construction
@@
As illustrated in Figure 2, the evidence library construction of the TrendFact follows a three-stage process: decomposition, generation, and retrieval. Specifically, we first employ an LLM to extract three or more key elements from each claim. We generate new claims for elements and retrieve supporting evidence from websites, which is then combined with the gold evidence from the initial TrendFact dataset to form a comprehensive evidence library containing 366,634 entries. More detailed descriptions are provided in Appendix E.
Category Label Acc Explain Cons Score
F-D F Full Discrepancy 0.2
F-C F Consistency 0.4
T-CD T Content Divergence 0.6
T-PC T Partial Consistency 0.8
T-FC T Full Consistency 1.0
Table 2: Definition of ECS. Label Acc indicates the accuracy of the fact verification, and Explain Cons represents the consistency of the explanation generated by the system and the gold one that evaluated by LLM .
4.5 ECS and HCPI Metrics Definition
To bridge the gap of evaluating explanation generation consistency and HPA of fact-checking systems, we introduce ECS and HCPI, respectively.
ECS
As shown in Table 2, we define five categories to describe explanation consistency based on the consistency between the gold explanation and the generated explanation with system’s verification accuracy: (1) F-D: Incorrect verification with full discrepancy; (2) F-C: Incorrect verification but consistent explanation; (3) T-CD: Correct verification with conflict explanation; (4) T-PC: Correct verification but partial consistency; (5) T-FC: Correct verification with full consistency. Higher explanation quality corresponds to higher scores when the system’s verification is correct.
Figure 3: Overview of FactISR. The bottom-right section shows the reasoning process of FactISR, where "flame" represents the reflection probability calculated based on influence score, which continuously decays with the number of iterative reflections. The reasoning process terminates when the maximum number of reflections is exceeded or no reflection step is sampled.
HCPI
The sample attributes of TrendFact align well with the key requirements for evaluating a fact-checking system’s Hotspot Perception Ability (HPA), as discussed in Section 1. Based on this connection, we introduce the Hotspot Claim Perception Index (HCPI) to quantitatively assess HPA by considering the social influence contributions of different claim types: Support (SUP), Refute (REF), and Not Enough Information (NEI). Overall, we define HCPI as follows:
$\text{HCPI}=\frac{S_{\text{SUP}}+S_{\text{REF}}+S_{\text{NEF}}}{\sum\limits_{i\in\text{All Claims}}\text{s}_{i}}$ (2)
Here, the three types of claim are weighted by different penalty factors in their social influence score calculations (the detailed formulas are in Appendix F). $\text{s}_{i}$ denotes the social influence score of the $i$-th sample, calculated from the weighted contributions of four hotspot indicators (views, discussions, engagements, and posts), combined with the risk score evaluated by GPT4.1, defined as:
$\text{s}_{i}=r_{i}\cdot\sum_{x\in\{v_{i},d_{i},e_{i},p_{i}\}}w_{x}\cdot\log(1+x)$ (3)
In summary, higher HCPI values correspond to stronger HPA performance of the system.